This is a test of the new technique for building windows. Use them to display pictures in text.

H e l l o

London. Michælmas Term lately over, and the Lord Chancellor sitting in Lincoln’s Inn Hall. Implacable November weather. As much mud in the streets,

as if the waters had but newly retired from the  face of the earth, and it
would not be wonderful to meet a Megalosaurus,  forty feet long or so, wad-
dling like an elephantine lizard up Holborn Hill.  Smoke lowering down from
chimney-pots, making a soft black drizzle, with  flakes of soot in it as big
as full-grown snow-flakes — gone into mourning,  one might imagine, for
the death of the sun. Dogs, indistinguishable in  mire. Horses, scarcely
better; splashed to their very blinkers. Foot pas-  sengers, jostling one an-
other’s umbrellas, in a general infection of ill-  temper, and losing their

foot-hold at street-corners, where tens of thousands of other foot passengers habe been slipping and sliding since the day broke (if this day ever broke), adding new deposits to the crust upon crust of mud, sticking at those points tenaciously to the pavement, and accumulating at compound interest.

H e l l o

London. Michælmas Term lately over, and the Lord Chancellor sitting in Lincoln’s Inn Hall. Implacable November weather. As much mud in the streets, as if the waters had but newly retired from the face of the earth, and it would not be wonderful to meet a Megalosaurus, forty feet long or so, waddling like an elephantine lizard up Holborn Hill. Smoke lowering down from chimney-pots, making a soft black drizzle, with flakes of soot

in it as big as full-grown snow-flakes — gone into mourning, one might imagine, for the death of the sun. Dogs, indistinguishable in mire. Horses, scarcely better; splashed to their very blinkers. Foot passengers, jostling one another’s umbrellas, in a general infection of ill-temper, and losing their foot-hold at street-corners, where tens of thousands of other foot passengers habe been slipping and sliding since the day broke (if this day ever broke), adding new deposits to the crust upon crust of mud, sticking at those points tenaciously to the pavement, and accumulating at compound interest.

Hello

London. Michælmas Term lately over, and the Lord Chancellor sitting in Lincoln’s Inn Hall. Implacable November weather. As much mud in the streets, as if the waters had but newly retired from the face of the earth, and it would not be wonderful to meet a Megalosaurus, forty feet long or so, waddling like an elephantine lizard up Holborn Hill. Smoke lowering down from chimney-pots, making a soft black drizzle, with flakes of soot in it as big as full-grown snow-flakes — gone into mourning, one might imagine, for the death of the sun. Dogs, indistinguishable in mire. Horses, scarcely better; splashed to their very blinkers. Foot passengers, jostling one another’s umbrellas, in a general infection of ill-temper, and losing their foot-hold at street-corners, where tens of thousands of other foot passengers habe been slipping and sliding since the day broke (if this day ever broke), adding new deposits to the crust upon crust of mud, sticking at those points tenaciously to the pavement, and accumulating at compound interest.

I hope these three examples make clear, what possibilties you have and how these routines are used.